I'm having trouble figuring out the Brain Game question titled CD Sort.
Your friend has a collection of CD: 2/3 of the CD's have booklets, 1/2 of the CD's are singles, and 1/4 are singles with booklets. All but two of the CDs are eith singles or have booklets or both. How many CDs does your friend have?

Thanks for your help.

Answer by scott8148(6628)

let x=CDs

1/4 are singles with booklets ___ x/4

1/2 of the CD's are singles ___ x/2 ___ but x/4 are singles w/booklets, so x/4 are singles w/o booklets

2/3 of the CD's have booklets ___ 8x/12 ___ but 3x/12 are singles w/booklets, so 5x/12 have booklets but are not singles

2 are neither singles nor have booklets

(3x/12)+(3x/12)+(5x/12)+2=x ___ (11x/12)+2=x ___ 11x+24=12x